OpenStudy (anonymous):

What are the domain and range of the function y=2√3x+4-5? x≥ -4/3 ; y ≥ -5 x≥ 4/3 ; y ≥ -5 x≤ -4/3 ; y ≤ -5 x≥ -4/3 ; y ≥5

OpenStudy (anonymous):

a

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

conti ?

OpenStudy (anonymous):

yes. one more question. just hold on. it may take awhile

OpenStudy (anonymous):

ok, i waiting

OpenStudy (anonymous):

x≥ -4/3 ; y ≥ -5

OpenStudy (anonymous):

is correct

OpenStudy (anonymous):

Which function is shown on the graph below? y = √2x-6-1 y = √2x+6+1 y = √2x-6+1 y = √2x+6=1

OpenStudy (anonymous):

thanks for waiting!

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

i have no more questions! you have solved all of them! ;)

OpenStudy (anonymous):

ok, see u later